# SSP-GNN: Learning to Track via Bilevel Optimization thanks: Corresponding author

### v2

@@ IV-A Train and Test Scenarios
Our synthetic data is generated using Stone Soup  [^2] https://github.com/dstl/Stone-Soup/, an open-source framework. We generate diverse scenarios with multiple targets using constant velocity and Ornstein-Uhlenbeck motion models. When generating detections, we fix the detection probability $(P_{D})$ at 1.0 except in Table LABEL:tab:p_d_exps which investigates the effects of an imperfect detector, and we fix measurement error standard deviation at $0.2$. We refer to target features as re-identification (ReID) as is common in computer vision. However, our approach can generalize to other features and sensing modalities, e.g., [15]. Our ReID features are drawn from a mixture of Gaussians (MoG) distributions. We control the “strength” of ReID features by changing the amount of overlap between the foreground and background distributions and define categories in terms of KL divergence: very weak (0.02 nats/0.029 bits), weak (0.5 nats/0.721 bits), moderate (3.125 nats/4.508 bits), and strong (12.5 nats/18.034 bits).
@@ IV-B Implementation Details
Given a pair of detections ${\bf y}_{i,t},{\bf y}_{i^{\prime},t^{\prime}}$ with $t\neq t^{\prime}$ we follow [16] to encode their spatio-temporal constraint into our latent representation. We denote the kinematic (spatial) component of a detection as ${\bf y}^{\mathbf{(s)}}_{i,t}$, ${\bf y}^{\mathbf{(s)}}_{i^{\prime},t^{\prime}}$, the ReID component as ${\bf y}^{\mathbf{(id)}}_{i,t}$, ${\bf y}^{\mathbf{(id)}}_{i^{\prime},t^{\prime}}$. For a node $v$ in $\mathcal{G}$, we define ${\bf h}^{(0)}_{v}=[{\bf y}^{\mathbf{(s)}}_{i},{\bf y}^{\mathbf{(id)}}_{i}]$. For an edge $e$ between the two detections in $\mathcal{G}$, we define ${\bf h}^{(0)}_{e}=[t^{\prime}-t,{\bf y}^{\mathbf{(s)}}_{i^{\prime}}-{\bf y}^{\mathbf{(s)}}_{i},\|{\bf y}^{\mathbf{(id)}}_{i^{\prime}}-{\bf y}^{\mathbf{(id)}}_{i}\|]$. We use 2D scenarios and 2D ReID features, i.e., ${\bf y}^{\mathbf{(s)}}\in\mathbb{R}^{2}$ and ${\bf y}^{\mathbf{(id)}}\in\mathbb{R}^{2}$.
Tracking graph and inference We instantiate $f_{\boldsymbol{\theta}}$ and learn parameters ${\boldsymbol{\theta}}$ on the detection graph $\mathcal{G}^{\textrm{det}}$ as shown in Fig. 2(a). Then, $f_{\boldsymbol{\theta}}(\mathcal{G}^{\textrm{det}})$ assigns scalar costs to the edges in $\mathcal{G}^{\textrm{det}}$. We copy these edge costs from $\mathcal{G}^{\textrm{det}}$ to $\mathcal{G}^{\textrm{trk}}$ by associating each edge $e_{i,j}\in\mathcal{G}^{\textrm{det}}$ with the edge between the pair of twin nodes associated with $i$ and the pair of twin nodes associated with $j$ in $\mathcal{G}^{\textrm{trk}}$. We assign a cost of 0 to the edges within the twin nodes, shown as red arrows in Figure 2(b). This formulation allows us to use the GNN-computed $\{c_{e}\}_{e=1}^{|\mathcal{E}|}$ as edge costs for the SSP algorithm to optimize over $\mathcal{G}^{\textrm{trk}}$. Detection and tracking graphs are represented via NetworkX  [^3] https://networkx.org/ except for GNN learning and inference the detection graph is converted to a Deep Graph Library (DGL) [29] object.
Training We use the Adam optimizer [13] with a learning rate of 0.001. Fig. 3 demonstrates a typical loss curve across 20 re-initialized training trials on the same train set. In practice we observe that $\ell_{1}$ tends to dominate total loss $\mathcal{L}$ because (a) our Stage I bootstrapping guarantees that $\ell_{2}\approx 0$ at epoch 0 and (b) $\ell_{2}$ captures the more trivial problem of simply keeping the ground truth paths sufficiently negative in cost, while $\ell_{1}$ addresses the more difficult problem of predicting the ground truth paths over competing sets of paths.
@@ IV-C Evaluation Metrics
Our method is inspired by [5, 16], etc., which were developed for tracking objects in video sequences. These methods tend to be evaluated with respect to Multiple Object Tracking Accuracy (MOTA) [4]. While our method is applicable to tracking extended targets in video sequences, we generate point observations in our Stone Soup simulation. We therefore report the MOTA metrics calculated with respect to the point observations for all of our experiments. We also report GOSPA [25] and SIAP[28] metrics when comparing our model’s performance to a filter-based tracker provided by Stone Soup.
Following [4], MOTA is defined as:
@@ IV-E Quantitative Evaluations
Comparison to baselines We compare our SSP-GNN tracker to an Extended Kalman Filter (EKF) based multi-object tracker and Edge-Belief GNN (based on [5]) . The EKF tracker implementation is provided by the Stone Soup library. Although EKF only processes the kinematic component of measurements and is an online method (while ours is offline/batch-based), it establishes a reference point in our comparison.
The Edge-Belief GNN algorithm [5] uses the same MPN architecture as our SSP-GNN, but instead of predicting edge costs, Edge-Belief GNN individually estimates the probability of an edge in the detection graph being part of a track or not. This is accomplished by modifying $g_{e}^{\mathrm{readout}}(\cdot,\cdot)$ to output sigmoid rather than logits. During inference, edge beliefs are rounded up or down to yield binary predictions, and a follow-on edge traversal heuristic ensures that the computed paths are node-disjoint. We observed that the Edge-Belief GNN tends to produce short (spurious) tracks; we added a post processing step which removes tracks shorter than a set length in order to strengthen the baseline (empirically we found minimum path threshold $=3$ was optimal).
@@
Refer to captionRefer to caption
@@
Edge-Belief $20$ Moderate 0.656 $\pm$ 0.032
SSP-GNN $20$ Moderate 0.891 $\pm$ 0.034
Edge-Belief $20$ Weak 0.633 $\pm$ 0.018
SSP-GNN $20$ Weak 0.873 $\pm$ 0.023

### v3

@@ IV-A Train and Test Scenarios
Our synthetic data is generated using Stone Soup  [^2] https://github.com/dstl/Stone-Soup/, an open-source framework. We generate diverse scenarios with multiple targets using constant velocity and Ornstein-Uhlenbeck motion models. When generating detections, we fix the detection probability $(P_{D})$ at 1.0 except in Table LABEL:tab:p_d_exps which investigates the effects of an imperfect detector, and we fix measurement error standard deviation at $0.2$. We refer to target features as re-identification (ReID) as is common in computer vision. However, our approach can generalize to other features and sensing modalities, e.g., [lenoach_feature-aided_2021]. Our ReID features are drawn from a mixture of Gaussians (MoG) distributions. We control the “strength” of ReID features by changing the amount of overlap between the foreground and background distributions and define categories in terms of KL divergence: very weak (0.02 nats/0.029 bits), weak (0.5 nats/0.721 bits), moderate (3.125 nats/4.508 bits), and strong (12.5 nats/18.034 bits).
@@ IV-B Implementation Details
Given a pair of detections ${\bf y}_{i,t},{\bf y}_{i^{\prime},t^{\prime}}$ with $t\neq t^{\prime}$ we follow [li_learning_tracking_2022] to encode their spatio-temporal constraint into our latent representation. We denote the kinematic (spatial) component of a detection as ${\bf y}^{\mathbf{(s)}}_{i,t}$, ${\bf y}^{\mathbf{(s)}}_{i^{\prime},t^{\prime}}$, the ReID component as ${\bf y}^{\mathbf{(id)}}_{i,t}$, ${\bf y}^{\mathbf{(id)}}_{i^{\prime},t^{\prime}}$. For a node $v$ in $\mathcal{G}$, we define ${\bf h}^{(0)}_{v}=[{\bf y}^{\mathbf{(s)}}_{i},{\bf y}^{\mathbf{(id)}}_{i}$]. For an edge $e$ between the two detections in $\mathcal{G}$, we define ${\bf h}^{(0)}_{e}=[t^{\prime}-t,{\bf y}^{\mathbf{(s)}}_{i^{\prime}}-{\bf y}^{\mathbf{(s)}}_{i},\|{\bf y}^{\mathbf{(id)}}_{i^{\prime}}-{\bf y}^{\mathbf{(id)}}_{i}\|$]. We use 2D scenarios and 2D ReID features, i.e., ${\bf y}^{\mathbf{(s)}}\in\mathbb{R}^{2}$ and ${\bf y}^{\mathbf{(id)}}\in\mathbb{R}^{2}$.
Tracking graph and inference We instantiate $f_{\boldsymbol{\theta}}$ and learn parameters $\boldsymbol{\theta}$} on the detection graph $\mathcal{G}^{\textrm{det}}$ as shown in Fig. 2(a). Then, $f_{\boldsymbol{\theta}}(\mathcal{G}^{\textrm{det}})$ assigns scalar costs to the edges in $\mathcal{G}^{\textrm{det}}$. We copy these edge costs from $\mathcal{G}^{\textrm{det}}$ to $\mathcal{G}^{\textrm{trk}}$ by associating each edge $e_{i,j}\in\mathcal{G}^{\textrm{det}}$ with the edge between the pair of twin nodes associated with $i$ and the pair of twin nodes associated with $j$ in $\mathcal{G}^{\textrm{trk}}$. We assign a cost of 0 to the edges within the twin nodes, shown as red arrows in Figure 2(b). This formulation allows us to use the GNN-computed $\{c_{e}\}_{e=1}^{|\mathcal{E}|}$ as edge costs for the SSP algorithm to optimize over $\mathcal{G}^{\textrm{trk}}$. Detection and tracking graphs are represented via NetworkX  [^3] https://networkx.org/ except for GNN learning and inference the detection graph is converted to a Deep Graph Library (DGL) [wang2019dgl] object.
Training We use the Adam optimizer [kingma2017adam] with a learning rate of 0.001. Fig. 3 demonstrates a typical loss curve across 20 re-initialized training trials on the same train set. In practice we observe that $\ell_{1}$ tends to dominate total loss $\mathcal{L}$ because (a) our Stage I bootstrapping guarantees that $\ell_{2}\approx 0$ at epoch 0 and (b) $\ell_{2}$ captures the more trivial problem of simply keeping the ground truth paths sufficiently negative in cost, while $\ell_{1}$ addresses the more difficult problem of predicting the ground truth paths over competing sets of paths.
@@ IV-C Evaluation Metrics
Our method is inspired by [Braso_2020_CVPR, li_learning_tracking_2022], etc., which were developed for tracking objects in video sequences. These methods tend to be evaluated with respect to Multiple Object Tracking Accuracy (MOTA) [bernardin_clear_mot]. While our method is applicable to tracking extended targets in video sequences, we generate point observations in our Stone Soup simulation. We therefore report the MOTA metrics calculated with respect to the point observations for all of our experiments. We also report GOSPA [Rahmathullah_2017] and SIAP[siap_votruba2001] metrics when comparing our model’s performance to a filter-based tracker provided by Stone Soup.
Following [bernardin_clear_mot], MOTA is defined as:
@@ IV-E Quantitative Evaluations
Comparison to baselines We compare our SSP-GNN tracker to an Extended Kalman Filter (EKF) based multi-object tracker and Edge-Belief GNN (based on [Braso_2020_CVPR]) . The EKF tracker implementation is provided by the Stone Soup library. Although EKF only processes the kinematic component of measurements and is an online method (while ours is offline/batch-based), it establishes a reference point in our comparison.
The Edge-Belief GNN algorithm [Braso_2020_CVPR] uses the same MPN architecture as our SSP-GNN, but instead of predicting edge costs, Edge-Belief GNN individually estimates the probability of an edge in the detection graph being part of a track or not. This is accomplished by modifying $g_{e}^{\mathrm{readout}}(\cdot,\cdot)$ to output sigmoid rather than logits. During inference, edge beliefs are rounded up or down to yield binary predictions, and a follow-on edge traversal heuristic ensures that the computed paths are node-disjoint. We observed that the Edge-Belief GNN tends to produce short (spurious) tracks; we added a post processing step which removes tracks shorter than a set length in order to strengthen the baseline (empirically we found minimum path threshold =3=3 was optimal).
@@
[…]
$Refer to caption\endlxSVG@picture$
@@
Edge-Belief 2020 Moderate 0.656 $\pm$ 0.032
SSP-GNN 2020 Moderate 0.891 $\pm$ 0.034
Edge-Belief 2020 Weak 0.633 $\pm$ 0.018
SSP-GNN 2020 Weak 0.873 $\pm$ 0.023